Comparative Statements of Shareholders’ Equity for Locke Intertechnology Corporation were reported as follows for the fiscal years ending December 31, 2019, 2020, and 2021.

LOCKE INTERTECHNOLOGY CORPORATION
Statements of Shareholders' Equity
For the Years Ended Dec. 31, 2019, 2020, and 2021
($ in millions)
Preferred
Stock,
$10 par
Common
Stock,
$1 par
Additional
Paid-in
Capital
Retained
Earnings
Total
Shareholders'
Equity
Balance at January 1, 2019 $ 55 $ 495 $ 1,878 $ 2,428
Sale of preferred shares 10 470 480
Sale of common shares, 7/1 9 81 90
Cash dividend, preferred (1 ) (1 )
Cash dividend, common (16 ) (16 )
Net income 290 290
Balance at December 31, 2019 10 64 1,046 2,151 3,271
Retirement of common shares, 4/1 (4 ) (36 ) (20 ) (60 )
Cash dividend, preferred (1 ) (1 )
Cash dividend, common (20 ) (20 )
3-for-2 split effected in the form of a
common stock dividend, 8/12
30 (30 )
Net income 380 380
Balance at December 31, 2020 10 90 980 2,490 3,570
10% common stock dividend, 5/1 9 90 (99 )
Sale of common shares, 9/1 3 31 34
Cash dividend, preferred (2 ) (2 )
Cash dividend, common (22 ) (22 )
Net income 412 412
Balance at December 31, 2021 $ 10 $ 102 $ 1,101 $ 2,779 $ 3,992

   
Required:
Infer from the statements the events and transactions that affected Locke Intertechnology Corporation's shareholders' equity and compute earnings per share as it would have appeared on the income statements for the years ended December 31, 2019, 2020, and 2021. No potential common shares were outstanding during any of the periods shown. (Enter your answers in millions rounded to 1 decimal place (i.e., 5,500,000 should be entered as 5.5).)
Answer #1

Earnings per share (2019) = (net income – preferred dividends)/( shares at Jan. 1 + new shares) = (290-1)/(55+(9*6/12)) = $4.86

Earnings per share (2020) = (net income – preferred dividends)/( shares at Jan. 1 + new shares) = (380-1)/((64*(1.50)-(4*9/12*1.50)) = $4.14

Earnings per share (2021) = (net income – preferred dividends)/( shares at Jan. 1 – retired shares) = (380-1)/((64*(1.50)-(4*9/12*1.50)) = $4.14

Earnings per share (2021) = (net income – preferred dividends)/( shares at Jan. 1 – retired shares) = (412-2)/((90*(1.10)+(3*4/12)) = $4.10